OpenStudy (anonymous):

i don't get this at all. i'm supposed to graph these equations: b+u=18 and b(2)+u=28 and then find which point they cross at...or the 'solution to system of equations' ?help? how do you even graph them?

OpenStudy (anonymous):

What course is this?

OpenStudy (anonymous):

They are linear equations. graph them like x + y = 18 and 2x + y = 28

OpenStudy (anonymous):

but how do you graph them.?

OpenStudy (anonymous):

There are a few ways. You can take each one and turn them into slope intercept form and graph them. For instance. x + y = 18 y = x + 18 Now that we have them in intercept form, we have are first point, which is the y intercept point of (0,18) and from there we know there is a rise of 1 and a run of 1 because are slope, which is 1x. We can use this rise and run to create the line or our next point. For instance, our next point will be from (0,18) will be (1,19) because we rise and run by 1. You can also just start entering points into the equation y = x + 18

OpenStudy (anonymous):

wouldn't the x be negative.?

OpenStudy (anonymous):

Yeah I made a boo boo :-)

OpenStudy (anonymous):

okay then resume please.

OpenStudy (anonymous):

so that mean the the rise and run is -1

OpenStudy (anonymous):

If you want to enter points, you would do something like this Just use some small number for x and put it into the equation Lets say x = 2, x = 3 and x = 4 y = -x + 18 y = -2 + 18 y = 16 first point is (2, 16) do the same for the other numbers. You can even use -2 and - 3y = -x + 18 y = -(-2) + 18 y = 2 + 18 y = 20 second point is (-2, 20)

OpenStudy (anonymous):

You see?

OpenStudy (anonymous):

yea. so just add the slope to the x&y.

OpenStudy (anonymous):

turn them into slope intercept form like I did and then graph them by using rise and run or by user points like I did. Also note, when you put it into slope intercept form, you have one point to start off with, which is the y intercept points. On the first one it is (0,18) on the second equation it is (0,28), which is handy when you want to use the rise and run method.

OpenStudy (anonymous):

kay. thank you.